How many ergs are there in three Joules?

  1. 10,000,000 ergs b. 20,000,000 ergs c. 30,000,000 ergs d. 40,000,000 ergs
1
Expert's answer
2021-03-29T16:47:42-0400
"3\ J\cdot\dfrac{10^7\ Erg}{1\ J}=3\cdot10^7\ Erg."

Answer:

c. 30,000,000 ergs